OpenStudy (anonymous):

From the ground, a hot-air balloon went up at a speed of 375 feet per minute for 5 minutes. It then went down at a speed of 225 feet per minute for 6 minutes. How far is the hot-air balloon above the ground? 150 feet 525 feet 1,125 feet 1,875 feet

Mehek (mehek14):

it went up 375/min for 5 mins so 375 * 5 = 1875 down 225/min for 6 mins 225 * 6 = 1350 subtract 1875 - 1350 = @DeacontheVexHunter

Mehek (mehek14):

subtract that and tell me what you get